For the reaction, the value of K_p is 41.0 (at 400
Answer #1

Kp1 = 41 T = 400K

Kp2 = X T = 705

HRxn = -92.2 kJ = -92200 J

this is exothermic so high T will not faovurKp

ln(Kp2/Kp1) = -H/R*(1/T1-1/T2)

ln(Kp2/41) = -92200/8.314*(1/400 - 1/705)

Kp2 = 41*exp(-92200/8.314*(1/400 - 1/705)) = 0.00025337562